OpenStudy (anonymous):

Solve (?) 4 (x + 3 ) + 6 ( x + 7 ) = 16

OpenStudy (anonymous):

3.8

OpenStudy (anonymous):

4x +12 +6x +42 = 16 10x +54 = 16 10x = - 38 x = -3.8

jimthompson5910 (jim_thompson5910):

4(x+3)+6(x+7)=16 4x+12+6x+42=16 10x+54=16 10x=16-54 10x=-38 x=-38/10 x=-19/5

OpenStudy (anonymous):

-3.8****